\begin{thebibliography}{15}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Abdelfattah et~al.(2022)Abdelfattah, Zhang, Fouda, Wang, and
  Wang]{abdelfattah2022g2netpl}
Rabab Abdelfattah, Xin Zhang, Mostafa~M Fouda, Xiaofeng Wang, and Song Wang.
\newblock G2netpl: Generic game-theoretic network for partial-label image
  classification.
\newblock \emph{British Machine Vision Conference (BMVC)}, 2022.

\bibitem[Abdelfattah et~al.(2023)Abdelfattah, Zhang, Wu, Wu, Wang, and
  Wang]{abdelfattah2023plmcl}
Rabab Abdelfattah, Xin Zhang, Zhenyao Wu, Xinyi Wu, Xiaofeng Wang, and Song
  Wang.
\newblock Plmcl: Partial-label momentum curriculum learning for multi-label
  image classification.
\newblock In \emph{Computer Vision--ECCV 2022 Workshops: Tel Aviv, Israel,
  October 23--27, 2022, Proceedings, Part II}, pp.\  39--55. Springer, 2023.

\bibitem[Berg et~al.(2012)Berg, Berg, Daume, Dodge, Goyal, Han, Mensch,
  Mitchell, Sood, Stratos, et~al.]{berg2012understanding}
Alexander~C Berg, Tamara~L Berg, Hal Daume, Jesse Dodge, Amit Goyal, Xufeng
  Han, Alyssa Mensch, Margaret Mitchell, Aneesh Sood, Karl Stratos, et~al.
\newblock Understanding and predicting importance in images.
\newblock In \emph{2012 IEEE Conference on Computer Vision and Pattern
  Recognition}, pp.\  3562--3569. IEEE, 2012.

\bibitem[Cole et~al.(2021)Cole, Mac~Aodha, Lorieul, Perona, Morris, and
  Jojic]{cole2021multi}
Elijah Cole, Oisin Mac~Aodha, Titouan Lorieul, Pietro Perona, Dan Morris, and
  Nebojsa Jojic.
\newblock Multi-label learning from single positive labels.
\newblock In \emph{Proceedings of the IEEE/CVF Conference on Computer Vision
  and Pattern Recognition}, pp.\  933--942, 2021.

\bibitem[Ke et~al.(2022)Ke, Zhu, Li, Shu, Qiao, and Ren]{ke2022hyperspherical}
Bo~Ke, Yunquan Zhu, Mengtian Li, Xiujun Shu, Ruizhi Qiao, and Bo~Ren.
\newblock Hyperspherical learning in multi-label classification.
\newblock In \emph{Computer Vision--ECCV 2022: 17th European Conference, Tel
  Aviv, Israel, October 23--27, 2022, Proceedings, Part XXV}, pp.\  38--55.
  Springer, 2022.

\bibitem[Kim et~al.(2022)Kim, Kim, Akata, and Lee]{kim2022large}
Youngwook Kim, Jae~Myung Kim, Zeynep Akata, and Jungwoo Lee.
\newblock Large loss matters in weakly supervised multi-label classification.
\newblock In \emph{Proceedings of the IEEE/CVF Conference on Computer Vision
  and Pattern Recognition}, pp.\  14156--14165, 2022.

\bibitem[Lin et~al.(2014)Lin, Maire, Belongie, Hays, Perona, Ramanan,
  Doll{\'a}r, and Zitnick]{lin2014microsoft}
Tsung-Yi Lin, Michael Maire, Serge Belongie, James Hays, Pietro Perona, Deva
  Ramanan, Piotr Doll{\'a}r, and C~Lawrence Zitnick.
\newblock Microsoft coco: Common objects in context.
\newblock In \emph{Computer Vision--ECCV 2014: 13th European Conference,
  Zurich, Switzerland, September 6-12, 2014, Proceedings, Part V 13}, pp.\
  740--755. Springer, 2014.

\bibitem[Spain \& Perona(2011)Spain and Perona]{spain2011measuring}
Merrielle Spain and Pietro Perona.
\newblock Measuring and predicting object importance.
\newblock \emph{International Journal of Computer Vision}, 91\penalty0
  (1):\penalty0 59--76, 2011.

\bibitem[Tsoumakas \& Katakis(2007)Tsoumakas and Katakis]{tsoumakas2007multi}
Grigorios Tsoumakas and Ioannis Katakis.
\newblock Multi-label classification: An overview.
\newblock \emph{International Journal of Data Warehousing and Mining (IJDWM)},
  3\penalty0 (3):\penalty0 1--13, 2007.

\bibitem[Verelst et~al.(2023)Verelst, Rubenstein, Eichner, Tuytelaars, and
  Berman]{verelst2023spatial}
Thomas Verelst, Paul~K Rubenstein, Marcin Eichner, Tinne Tuytelaars, and Maxim
  Berman.
\newblock Spatial consistency loss for training multi-label classifiers from
  single-label annotations.
\newblock In \emph{Proceedings of the IEEE/CVF Winter Conference on
  Applications of Computer Vision}, pp.\  3879--3889, 2023.

\bibitem[Wang et~al.(2016)Wang, Yang, Mao, Huang, Huang, and Xu]{wang2016cnn}
Jiang Wang, Yi~Yang, Junhua Mao, Zhiheng Huang, Chang Huang, and Wei Xu.
\newblock Cnn-rnn: A unified framework for multi-label image classification.
\newblock In \emph{Proceedings of the IEEE conference on computer vision and
  pattern recognition}, pp.\  2285--2294, 2016.

\bibitem[Wei et~al.(2015)Wei, Xia, Lin, Huang, Ni, Dong, Zhao, and
  Yan]{wei2015hcp}
Yunchao Wei, Wei Xia, Min Lin, Junshi Huang, Bingbing Ni, Jian Dong, Yao Zhao,
  and Shuicheng Yan.
\newblock Hcp: A flexible cnn framework for multi-label image classification.
\newblock \emph{IEEE transactions on pattern analysis and machine
  intelligence}, 38\penalty0 (9):\penalty0 1901--1907, 2015.

\bibitem[Xu et~al.(2022)Xu, Qiao, Lv, Geng, and Zhang]{xu2022one}
Ning Xu, Congyu Qiao, Jiaqi Lv, Xin Geng, and Min-Ling Zhang.
\newblock One positive label is sufficient: Single-positive multi-label
  learning with label enhancement.
\newblock \emph{Advances in Neural Information Processing Systems}, 2022.

\bibitem[Zhang et~al.(2021)Zhang, Cheng, Huang, Wen, Feng, Li, and
  Guo]{zhang2021simple}
Youcai Zhang, Yuhao Cheng, Xinyu Huang, Fei Wen, Rui Feng, Yaqian Li, and
  Yandong Guo.
\newblock Simple and robust loss design for multi-label learning with missing
  labels.
\newblock \emph{arXiv preprint arXiv:2112.07368}, 2021.

\bibitem[Zhou et~al.(2022)Zhou, Chen, Wang, Chen, and
  Heng]{zhou2022acknowledging}
Donghao Zhou, Pengfei Chen, Qiong Wang, Guangyong Chen, and Pheng-Ann Heng.
\newblock Acknowledging the unknown for multi-label learning with single
  positive labels.
\newblock In \emph{Computer Vision--ECCV 2022: 17th European Conference, Tel
  Aviv, Israel, October 23--27, 2022, Proceedings, Part XXIV}, pp.\  423--440.
  Springer, 2022.

\end{thebibliography}

\documentclass{article} 
\usepackage{iclr2023_conference_tinypaper,times}

\usepackage{hyperref}
\usepackage{url}
\usepackage{graphicx}
\usepackage{bbm}
\usepackage{amssymb}
\usepackage{subcaption}
\usepackage{amsmath}
\usepackage[font=small,labelfont=bf]{caption}

\title{Understanding Label Bias in Single Positive Multi-Label Learning}

\author{Julio Arroyo, Pietro Perona \& Elijah Cole\\
Caltech
}

\iclrfinalcopy 
\begin{document}

\maketitle
\vspace{-5mm}
\begin{abstract}
Annotating data for multi-label classification is prohibitively expensive because every category of interest must be confirmed to be present or absent. Recent work on single positive multi-label (SPML) learning shows that it is possible to train effective multi-label classifiers using only one positive label per image. However, the standard benchmarks for SPML are derived from traditional multi-label classification datasets by retaining one positive label for each training example (chosen uniformly at random) and discarding all other labels. In realistic settings it is not likely that positive labels are chosen uniformly at random. This work introduces protocols for studying label bias in SPML and provides new empirical results. 
\end{abstract}

\section{Introduction}
\emph{Multi-label classification} aims to predict the presence or absence of every category from a collection of categories of interest~\citep{tsoumakas2007multi}. This is a generalization of the standard \emph{multi-class classification} setting, in which the categories are mutually exclusive. Multi-label classification is often more realistic than multi-class classification in domains like natural image classification where multiple categories tend to co-occur~\citep{wang2016cnn,wei2015hcp}. 

The primary obstacle to training multi-label classifiers is the expense of exhaustively annotating the presence and absence of every category of interest. To mitigate this problem, \cite{cole2021multi} introduce the setting of \emph{single positive multi-label} (SPML) classification, which aims to train effective multi-label image classifiers using only one positive label per image. 

The SPML benchmarks introduced in \citet{cole2021multi} are artificially generated from fully annotated multi-label classification datasets. For each training image, one positive label is chosen uniformly at random to be retained and all other labels are discarded. While this simple model is a reasonable starting point, it is unlikely to reflect realistic annotator behavior~\citep{spain2011measuring}. If we want to apply SPML in realistic settings, we need to understand how different positive label selection biases affect the performance of SPML algorithms. 

Our key contributions are as follows: (1) We extend the SPML benchmarks in~\cite{cole2021multi} by generating new label sets based on different models of annotator bias. (2) We provide the first study of the effect of label bias in SPML, filling a crucial gap in the literature with significant implications for the real-world applicability of SPML. The code and data are publicly available.\footnote{\scriptsize\url{https://github.com/elijahcole/single-positive-multi-label/tree/spml-bias}}

\section{Related Work}

\textbf{SPML.} \citet{cole2021multi} formalized the SPML setting and introduced benchmarks and algorithms for the problem. Subsequent work has used these benchmarks to develop new SPML algorithms~\citep{kim2022large,abdelfattah2023plmcl,abdelfattah2022g2netpl,ke2022hyperspherical,zhou2022acknowledging,xu2022one,verelst2023spatial}. We complement this work by extending the benchmarks in~\citet{cole2021multi} and clarifying the role of label selection bias in SPML. 

\textbf{Models of annotator behavior.} The benchmarks in \citet{cole2021multi} are  unlikely to reflect the behavior of real annotators, who are more likely to mention an object if it is larger or closer to the center of the frame~\citep{berg2012understanding} or may neglect to mention it if it is ``too obvious"~\citep{spain2011measuring}. We take inspiration from this literature for our models of annotator behavior. 

\section{Bias Models for Single Positive Multi-Label Annotation}
Let $(\mathbf{x}, \mathbf{y})$ denote a training example, where $\mathbf{x} \in \mathbb{R}^p$ is an image and $\mathbf{y} \in \{0, 1\}^L$ is a label vector. Suppose we ask an annotator to name one class in $\mathbf{x}$. We want to model the probability $P(i)$ that class $i \in \{1, \ldots, L\}$ will be selected. As in~\cite{cole2021multi}, we assume $y_i = 0 \implies P(i) = 0$.

\noindent
\textbf{Uniform bias.} Choose one positive label uniformly at random: 
\begin{align}
    P_\mathrm{uniform}(i) &= \mathbbm{1}_{[y_i = 1]} \times 1/|\{i \in \{1, \ldots, L\} : y_i = 1\}|.\label{eq:uniform}
\end{align}

\noindent
\textbf{Size bias.} Choose a category in proportion to the fraction of the image it occupies:
\begin{align}
    P_\mathrm{size}(i) &= \mathbbm{1}_{[y_i = 1]} \times A_i / \sum_{j=1}^{L} A_j\label{eq:size}
\end{align}
where $A_i$ is the the sum of the bounding box areas of all instances of class $i$.

\noindent
\textbf{Location bias.} Choose a category more often if it occurs close to the center of the image:
\begin{align}
    P_\mathrm{location}(i) = \mathbbm{1}_{[y_i = 1]} \times D_i^{-1} / \sum_{j=1}^{L} D_j^{-1}\label{eq:location}
\end{align}
where $D_i$ is the average distance from the center of all instances of class $i$.

\noindent
\textbf{Semantic bias.} Determine probabilities based on empirical object spotting data: 
\begin{align}
    P_\mathrm{semantic}(i) &= \mathbbm{1}_{[y_i = 1]} \times f_i / \sum_{j=1}^L \mathbbm{1}_{[y_j = 1]} f_j\label{eq:semantic}
\end{align}
where $f_i$ is the empirical spotting frequency for category $i$. See Appendix~\ref{appendix:spotting} for details. 

\section{Results \& Discussion}

All experiments are conducted on COCO~\citep{lin2014microsoft}.
For each bias, we generate three independent realizations of SPML labels (see Appendix~\ref{appendix:label_gen}). Then for each dataset realization and loss, we use the code from~\citet{cole2021multi} to train and evaluate SPML methods under the ``Linear" protocol -- see \citet{cole2021multi} for implementation details. We present the results in Table~\ref{main-results}. Additional bias implementation details in Appendix~\ref{appendix:bias}.

\begin{table}[htb!]
\centering
\caption{Test set mean average precision (MAP) results for different bias models and training losses (AN = ``assume negative'' loss, AN-LS = ``assume negative'' loss with label smoothing, ROLE = regularized online label estimation, EM = entropy maximization). Values for $P_\mathrm{uniform}$ are copied from their respective papers.}
\label{main-results}
\resizebox{0.7\linewidth}{!}{
\begin{tabular}{|l|c|c|c|c|}
\hline
& $P_\mathrm{uniform}$ (\ref{eq:uniform})  &\bf $P_\mathrm{size}$ (\ref{eq:size}) &\bf $P_\mathrm{location}$ (\ref{eq:location}) &\bf $P_\mathrm{semantic}$ (\ref{eq:semantic}) \\
\hline
$\mathcal{L}_{\text{AN}}$~\citep{cole2021multi} &62.3 &57.0 $\pm$ 0.1 & 61.0 $\pm$ 0.2 & 59.8 $\pm$ 0.4 \\
\hline
 $\mathcal{L}_{\text{AN-LS}}$~\citep{cole2021multi} & 64.8 & 56.7 $\pm$ 0.3 & 62.7 $\pm$ 0.1 & 59.8 $\pm$ 0.1 \\
\hline
$\mathcal{L}_{\text{ROLE}}$~\citep{cole2021multi} & 66.3 & 60.1 $\pm$ 0.1 & 66.4 $\pm$ 0.0 & \underline{66.4} $\pm$ 0.0 \\
\hline
$\mathcal{L}_{\text{EM}}$~\citep{zhang2021simple} & \underline{70.7} & \underline{61.2} $\pm$ 0.1 & \underline{68.4} $\pm$ 0.0 & 65.6 $\pm$ 0.4 \\
\hline
\end{tabular}
}
\end{table}

We make three observations. 
(i) \emph{The ranking of different methods is similar across bias types.} This provides some evidence that the standard $P_\mathrm{uniform}$ benchmarks are a reasonable surrogate for more realistic biases that might be encountered in practice. 
(ii) \emph{Performance often drops significantly when we change from $P_\mathrm{uniform}$ to other biases.} Averaging over losses, the drop is largest for $P_\mathrm{size}$ (-7.3 MAP) and smallest for $P_\mathrm{location}$ (-1.4 MAP). 
(iii) \emph{Different losses have different sensitivities to label bias.} $\mathcal{L}_\mathrm{AN-LS}$ (-5.1 MAP) and $\mathcal{L}_\mathrm{EM}$ (-5.7 MAP) lose more performance on average than $\mathcal{L}_\mathrm{AN}$ (-3.1 MAP) and $\mathcal{L}_\mathrm{ROLE}$ (-2.0 MAP). 

These observations suggest that SPML algorithm rankings are roughly stable under different label biases, but the performance under $P_\mathrm{uniform}$ tends to overestimate the performance under other biases. We hope that our work will facilitate future research on SPML under realistic conditions. 

\textbf{Acknowledgements.} We thank Tsung-Yi Lin for providing the COCO object spotting data. 

\textbf{URM Statement.} At least one key author meets the URM criteria of ICLR 2023 Tiny Papers Track.

\bibliography{main}
\bibliographystyle{iclr2023_conference_tinypaper}

\appendix

\section{Implementation Details}

\subsection{Bias models}
\label{appendix:bias}

In this section we provide a few details for the non-trivial bias models: $P_\mathrm{size}$, $P_\mathrm{location}$, and $P_\mathrm{semantic}$. The equations are reproduced here for convenience, but are identical to those in the main paper. 

\textbf{Size bias.} We define 
\begin{align*}
    P_\mathrm{size}(i) &= \mathbbm{1}_{[y_i = 1]} \times A_i / \sum_{j=1}^{L} A_j
\end{align*}
where $A_i$ is the the sum of the bounding box areas of all instances of class $i$. Note that we do not make any modifications to handle intersecting bounding boxes, so $\sum_{j=1}^L A_j$ may be larger than the image size. 

\textbf{Location bias.} We define 
\begin{align*}
    P_\mathrm{location}(i) = \mathbbm{1}_{[y_i = 1]} \times D_i^{-1} / \sum_{j=1}^{L} D_j^{-1}
\end{align*}
where $D_i$ is the average distance from the center of all instances of class $i$. $D_i$ is implemented as the Euclidean distance between the center of the image and the center of the object bounding box. We also add a small value $\epsilon > 0$ to each $D_i$ to prevent division by zero for objects in the exact center of the image. $D_i^{-1}$ is defined to be zero if category $i$ is absent. 

\textbf{Semantic bias.} We define 
\begin{align*}
    P_\mathrm{semantic}(i) &= \mathbbm{1}_{[y_i = 1]} \times f_i / \sum_{j=1}^L \mathbbm{1}_{[y_j = 1]} f_j
\end{align*}
where $f_i$ is the empirical spotting frequency for category $i$, which is derived from annotator studies by the COCO team~\citep{lin2014microsoft}. Further details can be found in Appendix~\ref{appendix:spotting}.

\subsection{Label Set Generation}
\label{appendix:label_gen}

Like~\cite{cole2021multi}, we generate SPML training data by sampling labels from a fully-labeled multi-label classification dataset as follows: 
\begin{enumerate}
    \item Choose a bias model $P \in \{P_\mathrm{uniform}, P_\mathrm{size}, P_\mathrm{location}, P_\mathrm{semantic}\}$.
    \item Fix a random seed. 
    \item For each training image: (i) randomly choose one positive label according to $P$ and (ii) mark all other labels as ``unknown".
\end{enumerate}
For this work, we generate 3 realizations for each bias model with different seeds. Once generated, these SPML training sets are static, i.e. training labels are not re-sampled during training. The validation and test sets are always fully labeled. 

\section{COCO Object Spotting Data}
\label{appendix:spotting}

To build $P_\mathrm{semantic}$, we use object spotting data collected according to the protocols in~\citet{lin2014microsoft}. The data consists of annotations of the form \texttt{(pixel\_x, pixel\_y, category\_id)} for images in the COCO 2014 official validation set. Note that this data was collected by the COCO authors in 2017, and is different from the original object spotting data used to prepare the COCO 2014 dataset. 

We use this data to compute empirical object spotting frequencies $f_i$ for $i\in \{1, \ldots, L\}$ as follows:
\begin{enumerate}
    \item For each category, remove all spotting instances that do not fall inside of a bounding box of the same category.
    \item Count the remaining spotting instances for category $i$ across all images, which we call $f_i$. 
\end{enumerate}
We then plug in the $f_i$ values to the definition of $P_\mathrm{semantic}(i)$.
\end{document}